Week of 191202

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Kate (WLCG, chair), Maarten (WLCG, ALICE), Renato (LHCb), Vincent (security), Ben (computing), Paul (storage), Borja, Alberto (both monitoring)
  • remote: Christoph (CMS), Di (TRIUMF), Jens (NDGF), Marcelo (CNAF), Dave M (FNAL)

Experiments round table:

  • ATLAS reports ( raw view) -
    • No attendance due to ongoing ATLAS Software & Computing week - Indico
    • Nothing to report

  • CMS reports ( raw view) -
    • Central CMS production system run into a problem over the weekend
      • Sites see presently less production job pressure
      • Investigations are ongoing

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity:
      • Smooth running at ~100K jobs: MC simulation, WG productions, user analysis.
      • Had already staged 2015,2016,2017 data for the re-stripping campaign, awaiting for physics groups validation
    • Issues:
      • no significant issues

Sites / Services round table:

  • ASGC: nc
  • BNL: nc
  • CNAF: The HTCondor CE bulk migration is actually programed to January, also possible decommissioning of LSF middle January
Marcelo reported issues with web Vidyo. A ticket to Vidyo support was advised.
  • EGI: nc
  • FNAL: NTR
  • IN2P3: dCache has been upgraded to 5.2 version on last week maintenance. Upgrade went smoothly.
  • JINR: NTR
  • KISTI: nc
  • KIT:
    • atlassrm-kit.gridka.de has been updated to version 5.2 last week Wednesday. cmssrm-kit.gridka.de is next this week and lhcbsrm-kit.gridka.de will follow on Dec 11th.
    • Follow-up on last week's issue regarding the CVMFS signature keys rpm package (GGUS:144205): EGI states that installing its package (cvmfs-config-egi in UMD-4) is the correct choice. It will deploy the EGI signature key and CVMFS mount configuration on client machines that will enable access to a CVMFS path where all other keys, in particular CERN and OSG, can be found.
  • NDGF:
    • Last week the headnodes was upgraded from dCache 6.0.0 to 6.0.3 (with the pools remaining at 6.0.0). Unfortunately we didn't immediately notice that Xrootd (at least for Alice) was broken "Run: [ERROR] Server responded with an error: [3012] Failed to open file (Protocol Xrootd-4.0:109.105.124.161:37042 is not supported [27])". Late Friday afternoon the headnodes where downgraded to 6.0.0 again. So beware: dCache 6.0.3 headnodes are not compatible with 6.0.0 pools.
  • NL-T1: nc
  • NRC-KI: nc
  • OSG: nc
  • PIC: nc
  • RAL: nc
  • TRIUMF: NTR

  • CERN computing services: NTR for this week. Electrical maintenance is foreseen in P8 for 9th-16th December, jobs might be refused there during this time
  • CERN storage services: CVMFS repositories using new signing keys OTG0053606
  • CERN databases:
    • NTR
  • GGUS:
    • Another maintenance is planned for Wed this week
      • Release notes
      • A downtime has been scheduled for 07:30-10:00 UTC
      • Test alarms will be submitted as usual
  • Monitoring:
    • Draft availability reports for November sent around
  • MW Officer:
  • Networks: NTR
  • Security: NTR

AOB:

  • T1 upgrades to dCache >= v5.2.x required e.g. by DOMA TPC:
    • BNL, FNAL, TRIUMF: when do you plan to do the upgrade?
    • others: how did the upgrade go?

  • FNAL is already advanced with prep tasks (storage node upgrade to SL7 is 75% done), actual dCache upgrade is expected mid-January
  • TRIUMF needs yet to schedule the upgrade, it's possible it will be in January
  • Maarten advised sites to report all issues to dCache developers forums.
Edit | Attach | Watch | Print version | History: r20 < r19 < r18 < r17 < r16 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r20 - 2019-12-02 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback